Seth is two and a half times older than Pete. The difference in their ages is 45 years. Howa old are they?
Answer by Earlsdon(6103)

Let S be Seth's age and P be Pete's age. From the problem description, you can write:
1) S = 2.5P "Seth is two and a half times older than Pete"
2) S-P = 45 "The difference in their ages is 45 years" Substitute equation 1) into equation 2) and solve for P.
3) (2.5P)-P = 45
1.5P = 45 Divide both sides by 1.5
P = 30
S = 2.5(30)
S = 75
Seth is 75 years old and Pete is 30 years old.